# Morphing Tokens Draw Strong Masked Image Models

### v2

@@ 4.2 Overall Pipeline of Dynamic Token Morphing
We present a more advanced token morphing method, Dynamic Token Morphing (DTM), here. DTM’s dynamic nature stems from its simultaneous exploration of multiple and diverse token morphing cases. This design is based on the insight that morphing numerous tokens enhances the denoising effect while morphing fewer tokens retains detailed token representations. Essentially, the diversity in the process would yield significant benefits. Since morphing a fixed number of tokens may benefit only one side, DTM balances these advantages by morphing a diverse range of tokens. Furthermore, to achieve diversified morphed tokens, we divide the morphing process into multiple processes to create multiple morphed tokens. To this end, DTM encompasses three key components: 1) dynamic scheduler for token counts, 2) token morphing via scheduler, and 3) aligning morphed tokens. The overall framework of DTM is described in Fig. 4.
4.2.1 Dynamic scheduler.
@@
DTM generates multiple morphed tokens to ensure diversity and an extensive range of token variations. To this end, we first sample the final number of morphed tokens to remain after token morphing and the iteration number from uniform distributions, where represents the minimum number of morphed tokens and denotes the maximum number of iteration for sampling. Then, we define a token count scheduler , a sequence of token numbers that determines the number of tokens to morph for each iteration dynamically. Rather than sampling a sequence of random numbers that satisfies , we simply divide by for constant counts:
4.2.2 Token morphing via dynamic scheduler.
@@
The token morphing function is dynamically changed based on randomly sampled and . While our DTM is universal to any token matching methods, we employ bipartite matching [24] to search for contextually similar tokens efficiently. Specifically, we split tokens into two groups, with each token in the first group matched to its closest cosine similarity counterpart in the second group.
Eventually, we obtain the token morphing matrix , a contextual relation among tokens, from the target token representations as follows:
@@
where indicates that the token will be aggregated to the token . In addition, every token should be assigned to a specific cluster, even in cases where it forms a single token cluster, and each token should retain its exclusive association with a single cluster (i.e., and ). For the repetition , token pairs are successively matched.
Based on the token morphing matrix obtained by , the representations for online and target morphed tokens are derived by:
@@
4.2.3 Aligning morphed tokens.
We formulate the objective function by accumulating DTM losses, which aligns the representations of the corresponding online and target morphed tokens derived by DTM. The DTM loss with sampled and is formulated as follows:
where is a distance function and is a number of tokens aggregated for the online and target morphed tokens. Here, we utilize to consider all tokens aggregated for the morphed tokens. Note that our DTM loss can be extended to token-wise or image-level losses when or , respectively. We adopt Cosine distance for the distance function in Eq. 4 for all the DTM losses. To further enhance the dynamism of our method, we concurrently apply multiple DTMs. Therefore, the final objective function is the summation of all DTM losses, which is defined as:
where denotes the total number of simultaneously employed DTMs.
@@
| DeiT [42] | ICML 2021 | - | - | - | Label | - | 81.8 | - | - |
| DeiT-III [43] | ECCV 2022 | - | - | - | Label | - | 83.8 | 84.2 | 49.3 |
| Cosub [44] | CVPR 2023 | - | - | - | Label | 81.5 | 84.2 | 85.3 | 49.3 |
@@
| MoCo v3 [12] | ICCV 2021 | 300 | 300 | 300 | Pixel | 81.7 | 83.2 | 84.1 | 47.3 |
| BEiT [5] | ICLR 2021 | 300 | 800 | 800 | DALL-E | 81.7 | 83.2 | 85.2 | 47.1 |
| DINO [7] | ICCV 2021 | 3200 | 1600 | - | Feature | 82.0 | 83.6 | - | 46.8 |
| SplistMask [16] | arXiv 2021 | 300 | 300 | - | Pixel+Feat | 81.5 | 83.6 | - | 46.8 |
| iBOT [56] | ICLR 2022 | 3200 | 1600 | 1000 | Feature | 82.3 | 84.0 | 84.8 | 50.0 |
| MAE [19] | CVPR 2022 | - | 1600 | 1600 | Pixel | - | 83.7 | 85.6 | 48.1 |
| SimMIM [52] | CVPR 2022 | - | 800 | - | Pixel | - | 83.8 | - | - |
| MaskFeat [47] | CVPR 2022 | - | 1600 | 1600 | Feature | - | 84.0 | 85.7 | - |
| ExtreMa [50] | arXiv 2022 | - | 300 | - | Feature | - | 83.7 | - | 47.9 |
| CAE [11] | arXiv 2022 | 300 | 1600 | 1600 | Pixel+Feat | 82.0 | 83.9 | 86.3 | 50.2 |
| CMAE [22] | arXiv 2022 | - | 1600 | 1600 | Pixel+Feat | - | 84.4 | - | 50.1 |
| FD-CLIP [49] | arXiv 2022 | - | 300 | - | CLIP B/16 | - | 84.9 | - | 52.8 |
| BEiT v2 [34] | arXiv 2022 | - | 300 | 300 | CLIP B/16 | - | 85.0 | 86.6 | 52.7 |
| data2vec [4] | ICML 2022 | - | 800 | 1600 | Feature | - | 84.2 | 86.6 | - |
| mc-BEiT [27] | ECCV 2022 | - | 800 | 800 | VQGAN | - | 84.1 | 85.6 | 47.0 |
| MVP [48] | ECCV 2022 | - | 300 | 300 | CLIP B/16 | - | 84.4 | 86.3 | 52.4 |
| SdAE [13] | ECCV 2022 | - | 300 | - | Pixel | - | 84.1 | - | 48.6 |
| MSN [3] | ECCV 2022 | - | 600 | - | Feature | - | 83.4 | - | - |
| BootMAE [14] | ECCV 2022 | - | 800 | 800 | Pixel+Feat | - | 84.2 | 85.9 | 49.1 |
| CAN [31] | arXiv 2022 | - | 1600 | 800 | Pixel | - | 83.6 | 84.7 | - |
| SemMAE [26] | NIPS 2022 | - | 800 | - | Pixel | - | 83.3 | - | 46.3 |
| ConMIM [54] | ICLR 2023 | 300 | 800 | 1600 | Dictionary | 79.8 | 83.7 | 85.5 | 46.0 |
| RC-MAE [54] | ICLR 2023 | 1600 | 1600 | 1600 | Pixel | 82.0 | 83.6 | 86.1 | - |
| MixedAE [9] | CVPR 2023 | - | 1600 | - | Pixel | - | 83.9 | - | 49.8 |
| SIM [41] | CVPR 2023 | - | 1600 | - | Feature | - | 83.8 | - | - |
| HPM [46] | CVPR 2023 | - | 800 | 800 | Pixel | - | 84.2 | 85.8 | 48.5 |
| MIRL [21] | NeurIPS 2023 | - | 300 | 300 | Pixel | - | 84.1 | 85.4 | |
| DeepMIM [38] | arXiv 2023 | - | 300 | - | CLIP B/16 | - | 84.8 | - | - |
| dBOT [29] | ICLR 2024 | - | 1600 | 1600 | Feature | - | 84.5 | 86.6 | 49.5 |
| CrossMAE [17] | arXiv 2024 | 800 | 800 | 800 | Pixel | 79.3 | 83.7 | 85.4 | |
@@ 5 Experiment
5.0.1 Pre-training setup.
We pre-train our framework and other SSL frameworks through DTM on ImageNet-1K. We train our framework with ViT-S/16, ViT-B/16, and ViT-L/16 for 300 epochs using AdamW with momentum (0.9, 0.98) and a batch size of 1024. We use a learning rate of with cosine decay and warmup 10 epochs. We employ the CLIP base models [35] with its visual projector as a target model across all scales of ViT. Block-wise masking is used with a ratio of 0.4 following [5, 34]. Cosine distance is used as a distance metric for the objective according to an ablation study in Appendix. For the hyperparameters of DTM, we used 2 for , 1 for , 147 for , and 14 for all . We also pre-train various SSL frameworks through our DTM with the same fundamental setups. More detailed setups for pre-training are detailed in Appendix. All models are pre-trained using 8 V100-32GB GPUs.
@@
| Method | ViT-S | ViT-B | ViT-L |
| MIM (Baseline) | 82.7 | 84.3 | 86.4 |
| DTM | 83.2 | 85.4 | 86.7 |
| Gain | (+0.5) | (+1.1) | (+0.3) |
5.0.2 Fine-tuning setup.
We fine-tune our pre-trained models on ImageNet-1K [39] by default following the standard protocol [19, 34]. Pre-trained ViT-S, ViT-B, and ViT-L are fine-tuned for 300, 100, and 50 epochs, respectively. Optimization is performed with AdamW using a weight decay of 0.05. We use a layer-wise learning rate decay of 0.6 for ViT-S and ViT-B and 0.8 for ViT-L. Learning rate is set to with a linear warmup for 10 epochs for ViT-S and ViT-B and 5 epochs for ViT-L. Additional fine-tuning details are listed in Appendix.
5.1 ImageNet-1K Classification
We compare the fine-tuning accuracy of our method with previous state-of-the-art self-supervised methods on ImageNet-1K [39] datasets. The comparisons include supervised learning and SSL methods with various supervision signals. When the target model is CLIP, we only compare models pre-trained with CLIP B/16 for 300 epochs for a fair comparison. Table 3 reports the fine-tuning accuracies of ViT-S/B/L backbones. We observe that our MIM pre-trained by DTM achieves 83.2%, 85.4%, and 86.7% top-1 accuracies with ViT-S/16, ViT-B/16, and ViT-L/16, respectively, which outperforms state-of-the-art performances across the scales. Specifically, our method surpasses MVP [48], DeepMIM [38], and BEiT v2 [34] by 1.0p, 0.6p, and 0.4p on ViT-B/16, respectively. Moreover, our method outperforms other methods that leverage diverse supervision, demonstrating our method’s superiority among self-supervised learning methods.
5.2 Effectiveness of Our Method
5.2.1 Applicability on SSL frameworks.
We apply DTM on BEiT v2, MAE with the CLIP teacher [35], and BYOL [18] to verify its applicability to various SSL frameworks. As shown in Table 4, our DTM successfully improves the fine-tuning accuracy of MAE + CLIP, BEiT v2, and BYOL by 0.5p, 0.2p, and 0.4p, which reveals its general applicability to various frameworks beyond our baseline.
5.2.2 Scalability.
We further study the scalability of DTM. We compare the fine-tuning accuracies of our method to its baseline, which is pre-trained without DTM. We pre-train ViT-S/16, ViT-B/16, and ViT-L/16 with a resolution of for each method and evaluate the performance gains in fine-tuning accuracies. As shown in Table 5, DTM improved 0.4p, 1.1p, and 0.3p for ViT-S/16, ViT-B/16, and ViT-L/16, respectively, verifying its general efficacy across the scales.
| Target models | Method | ViT-S | ViT-B | ViT-L |
| SLIP | Baseline | 81.8 | 84.0 | 85.7 |
| DTM | 82.1 (+0.3) | 84.5 (+0.5) | 86.1 (+0.4) |
5.2.3 Further Applicability on SLIP.
We apply DTM with the SLIP [32] to verify its applicability to various target models. We compare fine-tuning accuracies of the baseline and our method pre-trained by target patches from SLIP. Specifically, we generate target tokens using SLIP along with its projector. We pre-train ViT-B/16 with a resolution of for 300 epochs and fine-tune for 100 epochs on ImageNet-1K [39] As shown in Table 6, our DTM successfully improves the fine-tuning accuracy of the baselines pre-trained with SLIP by 0.5p, which reveals its general applicability beyond CLIP [35].
5.2.4 Efficiency.
To verify the efficiency of our method, we compared Table 7 shows that bipartite matching is efficient and effective, significantly boosting the accuracy (+1.1%p) with only a 1% speed loss. While K-Means [30] also exhibits considerable improvements, it significantly degrades the training speed. Layer-wise K-Means shows accelerate the K-Means method by aggregating tokens within layers at the cost of degraded representations, leading to lower accuracy.
@@ 5.3 ADE20K Semantic Segmentation
We further evaluate semantic segmentation performance on ADE20K [55] to verify the transferability of our pre-trained model. We follow the training and evaluation protocol [19]; the models are fine-tuned for 160K iterations using UperNet [51] with a batch size of 16 and a resolution of 512512. We initialize UperNet with our pre-trained ViT-B/16. Detailed hyperparameters for semantic segmentation fine-tuning can be found in Appendix. The first right column in Table 3 shows the mIoU performance comparison. Our method also outperforms the previous state-of-the-art results with a margin of 0.3p, validating its superiority over other SSL methods. This result signifies that our method effectively enhances discriminability for dense prediction tasks.

### v3

@@ 4.2 Overall Dynamic Token Morphing Pipeline
We present a more advanced token morphing method, Dynamic Token Morphing (DTM), here. DTM’s dynamic nature stems from its simultaneous exploration of multiple and diverse token morphing cases. This design is based on the insight that morphing numerous tokens enhances the denoising effect while morphing fewer tokens retains detailed token representations. Furthermore, to achieve diversified morphed tokens, we divide the morphing process into multiple processes to create multiple morphed tokens. DTM encompasses three key components: 1) dynamic scheduler for token counts, 2) token morphing via scheduler, and 3) aligning morphed tokens. The overall framework of DTM is described in Fig. 4.
@@
Dynamic scheduler. DTM generates multiple morphed tokens to ensure diversity and an extensive range of token variations. To this end, we first sample the final number of morphed tokens to remain after token morphing and the iteration number from uniform distributions, where represents the minimum number of morphed tokens and denotes the maximum number of iteration for sampling. Then, we define a token count scheduler , a sequence of token numbers that determines the number of tokens to morph for each iteration dynamically. Rather than sampling a sequence of random numbers that satisfies , we simply divide by for constant counts:
@@
Token morphing via dynamic scheduler. The token morphing function is dynamically changed based on randomly sampled and . While our DTM is universal to any token matching methods, we employ bipartite matching (Karp et al., 1990) to search for contextually similar tokens efficiently. Specifically, we split tokens into two groups, with each token in the first group matched to its closest cosine similarity counterpart in the second group. Eventually, we obtain the token morphing matrix , a contextual relation among tokens, from the target token representations as follows:
@@
where indicates that the token will be aggregated to the token . In addition, we let every token be assigned to a specific cluster, even in cases where it forms a single token cluster, and each token should retain its exclusive association with a single cluster (i.e., and ). For the repetition , token pairs are successively matched. Based on the token morphing matrix obtained by , the representations for online and target morphed tokens are derived by:
@@
Aligning morphed tokens. We formulate the objective function by accumulating DTM losses, which aligns the representations of the corresponding online and target morphed tokens derived by DTM. The DTM loss with sampled and is formulated as follows:
where is a distance function and is a number of tokens aggregated for the online and target morphed tokens. Here, we utilize to consider all tokens aggregated for the morphed tokens. The DTM loss can be extended to token-wise or image-level losses when or , respectively. We adopt Cosine distance for the distance function in eq. 4 for all the DTM losses. To further enhance the dynamic nature of our method, we apply multiple DTM losses, each derived from its corresponding morphing case. The final objective function is the summation of all DTM losses, which is defined as:
where denotes the total number of simultaneously employed DTM losses.
@@
| DeiT (Touvron et al., 2021) | ICML 2021 | - | - | - | Label | - | 81.8 | - | - |
| DeiT-III (Touvron et al., 2022) | ECCV 2022 | - | - | - | Label | - | 83.8 | 84.2 | 49.3 |
| Cosub (Touvron et al., 2023) | CVPR 2023 | - | - | - | Label | 81.5 | 84.2 | 85.3 | 49.3 |
| MaskSub (Heo et al., 2023) | arXiv 2023 | - | - | - | Label | 81.7 | 84.2 | 85.3 | 50.2 |
@@
| MoCo v3 (Chen et al., 2021) | ICCV 2021 | 300 | 300 | 300 | Pixel | 81.7 | 83.2 | 84.1 | 47.3 |
| BEiT (Bao et al., 2021) | ICLR 2021 | 300 | 800 | 800 | DALL-E | 81.7 | 83.2 | 85.2 | 47.1 |
| DINO (Caron et al., 2021) | ICCV 2021 | 3200 | 1600 | - | Feature | 82.0 | 83.6 | - | 46.8 |
| SplistMask (El-Nouby et al., 2021) | arXiv 2021 | 300 | 300 | - | Pixel+Feat | 81.5 | 83.6 | - | 46.8 |
| iBOT (Zhou et al., 2022) | ICLR 2022 | 3200 | 1600 | 1000 | Feature | 82.0 | 84.0 | 84.8 | 50.0 |
| MAE (He et al., 2022) | CVPR 2022 | - | 1600 | 1600 | Pixel | - | 83.7 | 85.6 | 48.1 |
| SimMIM (Xie et al., 2022) | CVPR 2022 | - | 800 | - | Pixel | - | 83.8 | - | - |
| MaskFeat (Wei et al., 2022a) | CVPR 2022 | - | 1600 | 1600 | Feature | - | 84.0 | 85.7 | - |
| ExtreMa (Wu et al., 2022) | arXiv 2022 | - | 300 | - | Feature | 81.8 | 83.7 | - | 47.9 |
| CAE (Chen et al., 2022a) | arXiv 2022 | 300 | 1600 | 1600 | Pixel+Feat | 82.0 | 83.9 | 86.3 | 50.2 |
| CMAE (Huang et al., 2022) | arXiv 2022 | - | 1600 | 1600 | Pixel+Feat | - | 84.4 | - | 50.1 |
| FD-CLIP (Wei et al., 2022c) | arXiv 2022 | - | 300 | - | CLIP B/16 | - | 84.9 | - | 52.8 |
| BEiT v2 (Peng et al., 2022) | arXiv 2022 | - | 300 | 300 | CLIP B/16 | - | 85.0 | 86.6 | 52.7 |
| data2vec (Baevski et al., 2022) | ICML 2022 | - | 800 | 1600 | Feature | - | 84.2 | 86.6 | - |
| mc-BEiT (Li et al., 2022b) | ECCV 2022 | - | 800 | 800 | VQGAN | - | 84.1 | 85.6 | 47.0 |
| MVP (Wei et al., 2022b) | ECCV 2022 | - | 300 | 300 | CLIP B/16 | - | 84.4 | 86.3 | 52.4 |
| SdAE (Chen et al., 2022b) | ECCV 2022 | - | 300 | - | Pixel | - | 84.1 | - | 48.6 |
| MSN (Assran et al., 2022) | ECCV 2022 | - | 600 | - | Feature | - | 83.4 | - | - |
| BootMAE (Dong et al., 2022) | ECCV 2022 | - | 800 | 800 | Pixel+Feat | - | 84.2 | 85.9 | 49.1 |
| CAN (Mishra et al., 2022) | arXiv 2022 | - | 1600 | 800 | Pixel | - | 83.6 | 84.7 | - |
| SemMAE (Li et al., 2022a) | NIPS 2022 | - | 800 | - | Pixel | - | 83.3 | - | 46.3 |
| ConMIM (Yi et al., 2023) | ICLR 2023 | 300 | 800 | 1600 | Dictionary | 82.0 | 83.7 | 85.5 | 46.0 |
| RC-MAE (Yi et al., 2023) | ICLR 2023 | 1600 | 1600 | 1600 | Pixel | 82.0 | 83.6 | 86.1 | - |
| MixedAE (Chen et al., 2023) | CVPR 2023 | - | 1600 | - | Pixel | - | 83.9 | - | 49.8 |
| SIM (Tao et al., 2023) | CVPR 2023 | - | 1600 | - | Feature | - | 83.8 | - | - |
| HPM (Wang et al., 2023) | CVPR 2023 | - | 800 | 800 | Pixel | - | 84.2 | 85.8 | 48.5 |
| MIRL (Huang et al., 2024) | NeurIPS 2023 | - | 300 | 300 | Pixel | - | 84.1 | 85.4 | |
| DeepMIM (Ren et al., 2023) | arXiv 2023 | - | 300 | - | CLIP B/16 | - | 84.8 | - | - |
| dBOT (Liu et al., 2024) | ICLR 2024 | - | 1600 | 1600 | Feature | - | 84.5 | 86.6 | 49.5 |
| LUT (Kim et al., 2024) | ECCV 2024 | 400 | 1600 | 1600 | Pixel | 82.0 | 84.2 | 86.0 | 49.5 |
| CrossMAE (Fu et al., 2024) | arXiv 2024 | 800 | 800 | 800 | Pixel | 79.3 | 83.7 | 85.4 | - |
@@ 5 Experiment
In this section, we compare the ImageNet-1K classification and ADE20K segmentation performances. We explore the strengths of DTM including applicability on SSL frameworks and other pre-trained target models (e.g., SLIP), and training efficiency. We analyze the importance of dynamics and faster convergence of DTM. In appendix, we provide additional visualized examples of spatial inconsistencies. We also investigate the transferability of DTM and the applicability of superpixel algorithms. Then, we conduct ablation studies on various configurations of DTM. Implementation details including pre-training and fine-tuning details are listed in Section E of the Appendix.
| Target models | Method | ViT-S | ViT-B | ViT-L |
| SLIP | Baseline | 81.8 | 84.0 | 85.7 |
| DTM | 82.1 (+0.3) | 84.5 (+0.5) | 86.1 (+0.4) |
@@
5.1 ImageNet-1K Classification
We compare the fine-tuning accuracy of our method with previous state-of-the-art self-supervised methods on ImageNet-1K (Russakovsky et al., 2015) datasets. The comparisons include supervised learning and SSL methods with various supervision signals. When the target model is CLIP, we only compare models pre-trained with CLIP B/16 for 300 epochs for a fair comparison. Table 3 reports the fine-tuning accuracies of ViT-S/B/L backbones. Our baseline simply employ negative cosine loss with a vanilla CLIP model as the target model. We observe that our MIM pre-trained by DTM achieves 83.2%, 85.4%, and 86.7% top-1 accuracies with ViT-S/16, ViT-B/16, and ViT-L/16, respectively, which outperforms state-of-the-art performances across the scales. Specifically, our method surpasses MVP (Wei et al., 2022b), DeepMIM (Ren et al., 2023), and BEiT v2 (Peng et al., 2022) by 1.0p, 0.6p, and 0.4p on ViT-B/16, respectively. Moreover, our method outperforms other methods that leverage diverse supervision, demonstrating our method’s superiority among self-supervised learning methods.
Additionally, we extend DTM’s pre-training to 800 epochs, which improves fine-tuning accuracy of 85.5% on IN-1K. This result highlights 1) DTM merits longer pre-trainings and 2) DTM trained for 800 epochs surpasses others trained for 1600 epochs such as BeiT v2 (Peng et al., 2022). We believe even longer pre-training of DTM would lead to further improvements.
5.2 Effectiveness of Our Method
On other targets. To confirm our method’s applicability, we apply DTM with the SLIP (Mu et al., 2021) to verify its effectiveness across different target models. We compare fine-tuning accuracies of the baseline and our method pre-trained by target patches from SLIP. Specifically, we generate target tokens using SLIP along with its projector. We pre-train ViT-B/16 with a resolution of for 300 epochs and fine-tune for 100 epochs on ImageNet-1K (Russakovsky et al., 2015). As shown in Table 4, our DTM successfully improves the fine-tuning accuracy of the baselines pre-trained with SLIP by 0.5p, which reveals its general applicability beyond CLIP (Radford et al., 2021).
Applicability on SSL frameworks. We apply DTM on BEiT v2, MAE with the CLIP teacher (Radford et al., 2021), and BYOL (Grill et al., 2020) to verify its applicability to various SSL frameworks. As shown in Table 5, our DTM successfully improves the fine-tuning accuracy of MAE + CLIP, BEiT v2, and BYOL by 0.5p, 0.2p, and 0.4p, which reveals its general applicability to various frameworks beyond our baseline.
Efficiency. Table 6 shows that bipartite matching is efficient and effective, significantly boosting the accuracy (+1.1%p) with only a 1% speed loss. While K-Means (Lloyd, 1982) also exhibits considerable improvements, it significantly degrades the training speed. Layer-wise K-Means shows accelerate the K-Means method by aggregating tokens within layers at the cost of degraded representations, leading to lower accuracy.
Transferability. We verify the improved transferability of our pre-trained model. We compare fine-tuning accuracies of the baseline and our proposed model on iNaturalist datasets (Van Horn et al., 2018), which are highly imbalanced with different numbers of images per class, and Fine-Grained Visual Classification (FGVC) datasets. Table A and Table B in the Appendix show our DTM loss significantly improves the baseline with large margins, which reveals enhanced transferability.
@@ 5.3 ADE20K Semantic Segmentation
We further evaluate semantic segmentation performance on ADE20K (Zhou et al., 2017) to verify the transferability of our pre-trained model. We follow the training and evaluation protocol (He et al., 2022); the models are fine-tuned for 160K iterations using UperNet (Xiao et al., 2018) with a batch size of 16 and a resolution of 512512. We initialize UperNet with our pre-trained ViT-B/16. Detailed hyperparameters for semantic segmentation fine-tuning can be found in Appendix. The first right column in Table 3 shows the mIoU performance comparison. Our method also outperforms the previous state-of-the-art results with a margin of 0.3p, validating its superiority over other SSL methods. This result signifies that our method effectively enhances discriminability for dense prediction tasks.
5.4 Analysis
Ablation study on the importance of dynamics. Table 7 presents an ablation study on the effectiveness of the dynamic mechanism on DTM. We employ ViT-B/16 (Dosovitskiy et al., 2021) with a resolution of . The models are pre-trained for 300 epochs and fine-tuned for 100 epochs on ImageNet-1K (Russakovsky et al., 2015). For token morphing without the dynamic mechanism, half of the total 196 tokens are aggregated for each image. As shown in Table 7, token morphing with dynamic mechanism significantly improves the baseline while performance degrades without the dynamic mechanism, demonstrating the importance of dynamic mechanism in DTM.
Faster convergence of DTM. We analyze the behaviors of models pre-trained by DTM and token-wise objectives in Fig. 5. All the approaches employ CLIP representations for the target. As shown in Fig. 5(a), the fine-tuning accuracies of DTM surpass the accuracy of the token-wise supervision approaches. Fig. 5(b) shows that the model pre-trained by DTM exhibits a lower fine-tuning loss than the baseline model, verifying better convergence. Finally, Fig. 5(c) verifies the consistent effect of DTM across various pre-training epochs.
| Method | Dynamic | FT () | |
| Baseline | - | - | 84.3 |
| DTM | ✓ | - | 84.0 (-0.3) |
| DTM | ✓ | ✓ | 85.4 (+1.1) |
6 Conclusion
Appendix
This appendix includes additional experimental analyses of our proposed method:
Section A: Additional examples of spatial inconsistency among patches from pre-trained models
Section B: Transferability of DTM on iNaturalist and Fine-Grained Visual Classification (FGVC) datasets
Section D: Ablation studies on the number of morphing schedules and effects of randomness in the number of morphing tokens, randomness in gradual token morphing, and target normalization
Appendix A More Examples on Spatial Inconsistency
We further investigate the spatial inconsistency of patch representations generated by a pre-trained model across various samples. We visualize the token-wise zero-shot classification results without and with token aggregation using CLIP (Radford et al., 2021). Specifically, correct and incorrect tokens are marked in green and red, respectively, with a gradient to darker shades of red, indicating a more significant deviation from the true class. . As shown in Fig. A, patch representations without token aggregation reveal spatially inconsistent patch-wise predictions compared to the prediction results with token aggregation, which reveals the spatially inconsistent representations among patches. In addition, predictions with token aggregation exhibit significantly more correctly predicted patches than predictions without token aggregation.
Appendix B Investigation on Transfer learning
iNaturalist datasets. We verify the improved transferability of our pre-trained model. We compare fine-tuning accuracies of the baseline and our proposed model on iNaturalist 2018, iNaturalist 2019, and mini iNaturalist 2021 (Van Horn et al., 2018), which are highly imbalanced with different numbers of images per class. All the models are ViT-B/16 with a resolution of 224224. Following the protocol (Kornblith et al., 2019), we perform grid searches on learning rates and weight decay and report the maximum accuracy and the mean and standard deviation of the accuracies. Table A shows our DTM loss significantly improves the baseline with large margins, which reveals enhanced transferability.
Fine-Grained Visual Classification (FGVC) datasets. We further validate the transferability of our method. Following the evaluation protocol as above, we conduct comparisons on FGVC datasets. Specifically, we evaluate fine-tuning accuracies on Birds, CUB-200, CIFAR-10, CIFAR-100, and Dogs through grid searches with different learning rates and weight decays. As shown in Table B, our method outperforms the baseline overall across the datasets, which shows superior transferability and tuning robustness.
Appendix C Compatibility with Superpixel Algorithms
We perform additional experiments to use the concept of superpixels (Achanta et al., 2012; Chang et al., 2023) in ours, both directly to tokens (layer-wise superpixel) and can be combined with DTM. The layer-wise superpixel method uses constant numbers of superpixels and iterations. Table C shows superpixel-based methods enjoy notable gains but do not exceed the bipartite matching one. While the layer-wise superpixel method also utilizes superpixels, the layer-wise token aggregation across the encoder layers bipartite matching approach risks harming intermediate representations during encoding
C.1 Ablation Study on the Number of Schedules
Our method can further enhance the diversity of target morphed tokens by employing multiple schedules within a single iteration. Thus, we study the effects of learning diverse morphed tokens derived from multiple morphing schedules simultaneously. We compare the fine-tuning accuracy of models pre-trained by our DTM, varying the number of schedules. We pre-train ViT-B/16 with a resolution of for 100 epochs and fine-tune for 100 epochs on ImageNet-1K (Russakovsky et al., 2015) and ImageNet-100 (Russakovsky et al., 2015). Table D reveals that exploring diverse target morphed tokens improves the representation capability of pre-trained models, leading to increased fine-tuning accuracies of at least 0.1p and 0.2p on ImageNet-1K and ImageNet-100 compared to the model pre-trained by the single DTM approach, respectively. However, the goal of experiencing diverse morphed tokens at once appears to be attained with double scheduling, resulting in no additional gains in performance through further exploration. Given that utilizing two schedules yields the best and is most efficient among all other multiple scheduling options, we adopt the double morphing scheduling approach.